OpenStudy (anonymous):

On the Fourth of July, 850 soft drinks were sold at a concession stand, some at $.72 each, the others at $.90 each. The receipts were $690. How many were sold at each price. Use a system!

OpenStudy (anonymous):

x+y=850 .72()+.90()=690

OpenStudy (anonymous):

Thks

OpenStudy (anonymous):

Can someone solve it?

OpenStudy (amistre64):

c + p = 850 c = 850 - p ........................ c(.72) + p(.90) = 690 (850 - p)(.72) + p(.90) = 690 850(.72) -.72p +.90p = 690 p(.90 -.72) = 690 - 850(.72) p = 690 - 850(.72) ----------- .90 -.72 p = 433.333333

OpenStudy (anonymous):

\[.72x+.9(850-x)=690\] should work

OpenStudy (amistre64):

kinda odd to sell 1/3 of a pepsi lol

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

that is about as much as i could drink. i am getting cranky time for lunch

OpenStudy (amistre64):

as i see it, someone was skiming off the profits

OpenStudy (anonymous):

two thirds pepsi, one third rum maybe?

OpenStudy (amistre64):

:) i may be a bad mormon in somethings; but that aint one of them lol

OpenStudy (anonymous):

;)

OpenStudy (amistre64):

it might bode better to switch the coke and pepsi prices and see if a discrete amount comes up